How much would $500 invested at 8% interest compounded annually be worth after 4 years? Round your answer to the nearest cent. Do not include units in your answer. A(t)=P(1+r/n)(r*t)
Answer by stanbon(75887)

How much would $500 invested at 8% interest compounded annually be worth after 4 years? Round your answer to the nearest cent. Do not include units in your answer. A(t)=P(1+(r/n))^(n*t)
----------------------
A(4) = 500(1+(0.08/1))^(1*4)
----
A(4) = 500(1.08)^4
----
A(4) = $680.24
==================
Cheers,
Stan H.